# Photometric White Dwarf Rotation

### v1

@@ 5 Conclusion
144002497 | 7.6213 0.0022 | 48.042 | 1.00 | 42.8 | 72910 | 0.606 | 2/2 | |
837076117 | 11.468 0.007 | 7.899 | 1.00 | 11.2 | 81335 | 0.65 | 1/2 | DA |
307982318 | 1.53006 7e-05 | 4.803 | 1.00 | 5.0 | 31880 | 0.67 | 1/3 | hotDAV |
88564975 | 8.9702 0.003 | 173.775 | 1.00 | 92.9 | 80000 | 0.67 | 3/3 | |
136884288 | 0.6053 0.0006 | 7.730 | 1.00 | 13.6 | 8440 | 0.81 | 1/1 | DAHe |
383647517 | 13.924 0.014 | 13.263 | 1.00 | 6.2 | 6754 | 0.545 | 7/16 | DQpec |
1102584065 | 26.187 0.022 | 17.910 | 1.00 | 9.1 | 110000 | 0.57 | 12/13 | DO |
436747174 | 4.166 0.026 | 290.193 | 1.00 | 59.7 | 22500 | 0.532 | 1/1 | DA-CV+M |
434196824 | 3.904 0.023 | 53.066 | 1.00 | 62.4 | 14491 | 0.258 | 1/1 | |
345265261 | 3.759 0.022 | 7.601 | 1.00 | 12.2 | 38780 | 0.843 | 1/1 | DA |
321272216 | 4.408 0.031 | 34.385 | 0.99 | 40.0 | 12597 | 0.117 | 1/1 | |
248861930 | 12.93 0.26 | 23.353 | 1.00 | 28.0 | 36903 | 0.521 | 1/1 | DBA |
247992671 | 15.4 0.4 | 8.742 | 1.00 | 10.8 | 8000 | 0.23 | 1/1 | |
241256067 | 15.778 0.028 | 5.352 | 1.00 | 6.0 | 4730 | 0.495 | 1/2 | DC |
124894242 | 0.32555 0.00017 | 4.075 | 1.00 | 6.9 | 15919 | 0.435 | 1/1 | |
154942231 | 4.4236 0.0022 | 7.839 | 1.00 | 6.4 | 12000 | 0.15 | 3/5 | |
649932200 | 0.9712 0.0005 | 31.107 | 1.00 | 28.4 | 26461 | 1.036 | 3/3 | |
52584399 | 3.801 0.008 | 112.873 | 1.00 | 66.8 | 19505 | 0.249 | 3/3 | |
1883129777 | 1.9446 0.0004 | 143.936 | 1.00 | 71.9 | 16480 | 0.229 | 3/3 | |
1881564044 | 6.137 0.004 | 6.042 | 1.00 | 7.4 | 34000 | 0.2 | 2/3 | DAH |
1101460688 | 23.24 0.03 | 10.905 | 1.00 | 6.9 | 88000 | 0.656 | 2/3 | |
22014729 | 7.6477 0.0032 | 135.408 | 1.00 | 55.8 | 36580 | 0.44 | 1/2 | |
31798074 | 21.828 0.017 | 42.369 | 1.00 | 5.3 | 12011 | 0.398 | 1/10 | |
262921847 | 13.892 0.004 | 52.290 | 1.00 | 20.7 | 7000 | 0.23 | 12/12 | |
311886800 | 10.2157 0.0025 | 32.576 | 1.00 | 19.4 | 100000 | 0.6 | 6/6 | |
178994042 | 0.2138 7e-05 | 4.400 | 1.00 | 5.6 | 19770 | 0.57 | 1/1 | |
220478684 | 16.611 0.010 | 25.143 | 1.00 | 12.6 | 44589 | 0.483 | 19/20 | |
381976323 | 8.722 0.01 | 133.547 | 1.00 | 45.2 | 4000 | 0.2 | 11/11 | |
235072034 | 84.0 4.0 | 20.540 | 1.00 | 10.5 | 3500 | 0.23 | 2/3 | |
219244444 | 7.2886 0.0027 | 245.106 | 1.00 | 12.5 | 6440 | 0.25 | 3/5 | DA7+M+planet |
245834374 | 33.6 1.8 | 7.275 | 1.00 | 6.8 | 34210 | 0.69 | 1/1 | DA+M |
85146644 | 5.317 0.0011 | 68.977 | 1.00 | 51.4 | 60260 | 0.41 | 3/3 | |
107310223 | 11.28 0.23 | 21.505 | 1.00 | 30.9 | 48116 | 0.366 | 1/1 | DA1+dM |
1101900758 | 2.65191 0.0003 | 31.335 | 1.00 | 21.4 | 23941 | 0.398 | 5/5 | |
160119816 | 21.453 0.017 | 112.794 | 1.00 | 43.3 | 90000 | 0.56 | 9/9 | |
300284179 | 12.65 0.29 | 12.426 | 1.00 | 13.8 | 38490 | 0.54 | 1/1 | |
126910998 | 19.4 0.6 | 36.753 | 1.00 | 54.9 | 23075 | 0.819 | 1/1 | DA+M |
218971976 | 6.7081 0.001 | 263.431 | 1.00 | 65.1 | 39260 | 0.38 | 2/2 | DA+M |
321979116 | 2.69506 0.00016 | 92.402 | 1.00 | 36.5 | 8650 | 0.96 | 2/2 | DAP5.8 |
183533908 | 2.55267 0.00024 | 11.122 | 1.00 | 13.6 | 7638 | 0.74 | 1/1 | |
630539472 | 1.592 0.004 | 4.231 | 1.00 | 7.4 | 10027 | 0.821 | 1/1 | |
201255204 | 6.34 0.03 | 59.895 | 1.00 | 52.0 | 4000 | 0.2 | 2/2 | DBH |
439917321 | 12.98 0.28 | 61.269 | 1.00 | 48.6 | 19193 | 0.57 | 1/1 | DA+M |
55096188 | 2.75096 0.00021 | 4.379 | 1.00 | 6.5 | 10000 | 0.6 | 2/3 | |
53190694 | 9.923 0.004 | 12.529 | 1.00 | 9.3 | 30502 | 0.41 | 3/4 | DA3+dM |
952264940 | 1.78894 0.00018 | 7.713 | 1.00 | 8.6 | 25692 | 1.151 | 1/2 | DBH |
219769847 | 8.72 0.13 | 13.789 | 1.00 | 20.7 | 21620 | 0.72 | 1/1 | DA3+dM |
292114384 | 0.29255 0.00014 | 4.110 | 1.00 | 5.1 | 18943 | 0.57 | 1/1 | |
66714896 | 12.579 0.009 | 10.848 | 1.00 | 8.8 | 63073 | 0.69 | 1/3 | |
349233389 | 52.0 5.0 | 31.595 | 1.00 | 25.7 | 23690 | 0.58 | 1/1 | DA2+M |
335501964 | 6.7654 0.0025 | 5.380 | 1.00 | 8.1 | 90000 | 0.53 | 1/2 | |
165916724 | 0.092325 5e-07 | 4.156 | 1.00 | 5.2 | 9557 | 0.883 | 1/3 | DAH |
274039489 | 7.32 0.09 | 67.842 | 1.00 | 58.5 | 42370 | 0.66 | 1/1 | DA1.2+pair |
328029653 | 14.157 0.004 | 19.335 | 0.99 | 12.5 | 6340 | 0.53 | 5/5 | |
311920919 | 0.1614889 1e-06 | 4.611 | 1.00 | 6.2 | 14116 | 0.68 | 1/4 | |
155871645 | 0.1385659 8e-07 | 6.552 | 1.00 | 5.5 | 34050 | 0.77 | 3/5 | |
344927121 | 2.53418 0.00018 | 113.339 | 1.00 | 67.3 | 26088 | 0.381 | 4/4 | |
148719841 | 8.549 0.004 | 17.920 | 1.00 | 18.7 | 6169 | 0.72 | 2/2 | |
415714190 | 0.384926 4e-06 | 9.333 | 1.00 | 8.4 | 7690 | 0.636 | 3/3 | |
277713491 | 8.9986 0.0032 | 11.795 | 1.00 | 9.2 | 7518 | 0.578 | 3/4 | |
471013515 | 6.05 0.07 | 27.262 | 1.00 | 34.8 | 55000 | 0.6 | 1/1 | |
5393020 | 1.5786 0.00014 | 54.426 | 1.00 | 49.4 | 8000 | 0.35 | 1/2 | |
610754879 | 0.91146 9e-05 | 11.175 | 1.00 | 10.9 | 47617 | 1.341 | 2/3 | |
365207316 | 2.539 0.005 | 6.316 | 1.00 | 7.1 | 10520 | 0.226 | 2/2 | |
425079955 | 17.076 0.016 | 20.821 | 1.00 | 16.2 | 14252 | 0.285 | 2/2 | |
1271382181 | 7.426 0.005 | 40.083 | 1.00 | 21.8 | 24015 | 0.25 | 8/8 | |
115970839 | 2.6994 0.0004 | 41.932 | 1.00 | 30.0 | 8825 | 0.219 | 5/5 | DA+M: |
159819681 | 10.728 0.006 | 29.967 | 1.00 | 22.8 | 27876 | 0.573 | 5/5 | |
158971558 | 6.327 0.004 | 7.300 | 1.00 | 7.3 | 14319 | 0.323 | 2/2 | |
1550807820 | 0.566897 1.6e-05 | 5.636 | 1.00 | 8.6 | 44523 | 1.293 | 5/5 | |
7622101 | 8.5708 0.0028 | 11.742 | 1.00 | 7.6 | 14159 | 0.2 | 7/7 | |
349993999 | 4.381 0.0005 | 7.642 | 1.00 | 7.0 | 24751 | 0.756 | 1/1 | |
1688534307 | 0.0902002 4e-07 | 4.253 | 1.00 | 5.3 | 66522 | 0.637 | 1/2 | |
62918265 | 4.6612 0.0008 | 12.845 | 0.99 | 13.1 | 8133 | 0.724 | 6/6 | |
1922433415 | 2.03387 0.00015 | 129.416 | 1.00 | 60.4 | 21155 | 0.214 | 2/2 | |
100469915 | 12.936 0.006 | 18.310 | 1.00 | 21.6 | 12644 | 0.152 | 2/2 | |
2026050241 | 8.0844 0.0024 | 6.393 | 1.00 | 5.7 | 48831 | 0.489 | 1/2 | |
60112260 | 7.37 0.1 | 14.436 | 1.00 | 20.0 | 7682 | 0.756 | 1/1 | |
269937578 | 7.2969 0.0019 | 79.671 | 1.00 | 45.4 | 32755 | 0.274 | 4/5 | |
2052430471 | 0.125404 2.9e-05 | 4.277 | 1.00 | 5.4 | 28678 | 0.368 | 1/1 | |
124573902 | 7.9918 0.0035 | 21.546 | 1.00 | 21.9 | 12173 | 0.371 | 2/2 | wd+pair |
270432557 | 9.311 0.01 | 6.316 | 1.00 | 5.4 | 10000 | 0.23 | 2/3 | |
43529091 | 9.44 0.14 | 11.479 | 1.00 | 14.0 | 11006 | 0.17 | 1/1 | |
278861557 | 3.1624 0.0004 | 26.520 | 1.00 | 7.8 | 10460 | 0.23 | 19/20 | wd+pair |
62846288 | 0.500906 1.4e-05 | 23.831 | 1.00 | 17.9 | 16268 | 0.935 | 3/4 | |
56813164 | 8.05 0.11 | 25.988 | 1.00 | 34.7 | 17304 | 0.39 | 1/1 | DA+M |
953086708 | 0.088138 1.3e-05 | 5.293 | 1.00 | 7.4 | 7735 | 0.564 | 1/1 | DAEH |
95475692 | 1.94319 0.00021 | 45.117 | 1.00 | 36.2 | 13737 | 0.428 | 2/2 | DA+M |
68015843 | 3.015 0.0018 | 21.609 | 1.00 | 21.7 | 12392 | 0.4 | 2/2 | DA+M4.5 |
219485299 | 1.9770 0.0020 | 4.513 | 1.00 | 18.9 | 10307 | 0.63 | 3/3 | DA+M |
1507548590 | 1.9272 0.0004 | 50.124 | 1.00 | 49.2 | 25089 | 0.214 | 2/2 | |
1101327387 | 1.30545 9e-05 | 21.739 | 1.00 | 19.5 | 15659 | 1.13 | 2/2 | DAH |
416538823 | 16.021 0.011 | 335.044 | 1.00 | 26.1 | 18840 | 0.8 | 4/4 | DA+M |
7983187 | 14.32 0.16 | 77.693 | 1.00 | 45.2 | 17210 | 0.59 | 2/2 | DA3+dM |
251857373 | 2.31675 0.00012 | 175.807 | 1.00 | 110.6 | 40698 | 0.56 | 2/2 | |
408015814 | 29.52 0.05 | 13.089 | 1.00 | 11.8 | 76010 | 0.061 | 2/2 | |
200725303 | 9.48 0.07 | 35.024 | 1.00 | 33.5 | 90000 | 0.75 | 2/2 | |
159894126 | 0.46415 0.00017 | 4.410 | 1.00 | 5.9 | 8627 | 0.613 | 1/2 | |
651923454 | 42.4 2.9 | 16.656 | 1.00 | 11.9 | 21644 | 0.746 | 1/1 | wd+pair |
651337335 | 30.5 1.5 | 6.331 | 1.00 | 9.7 | 21796 | 0.217 | 1/1 | |
650999219 | 24.6 0.5 | 24.641 | 1.00 | 26.3 | 70000 | 0.49 | 2/2 | |
93031595 | 9.94 0.16 | 17.306 | 1.00 | 25.7 | 8751 | 0.911 | 1/1 | DZ |
685012668 | 2.8735 0.0018 | 107.688 | 1.00 | 44.9 | 20041 | 0.238 | 5/5 | |
672383331 | 17.8 0.5 | 7.107 | 1.00 | 10.9 | 21076 | 0.279 | 1/1 | |
257689590 | 13.35 0.13 | 5.132 | 0.99 | 5.2 | 8629 | 0.183 | 2/2 | |
455072135 | 11.36 0.21 | 90.850 | 1.00 | 49.0 | 12965 | 0.116 | 1/1 | EB |
941411508 | 10.547 0.006 | 8.630 | 1.00 | 9.1 | 40000 | 0.556 | 2/2 | |
942842236 | 8.233 0.004 | 10.210 | 1.00 | 10.1 | 77164 | 0.75 | 2/2 | |
332828037 | 2.07518 0.00024 | 15.973 | 1.00 | 17.4 | 13936 | 0.382 | 2/2 | |
939953390 | 29.85 0.05 | 9.949 | 1.00 | 6.7 | 63208 | 0.667 | 3/3 | |
952000432 | 21.158 0.025 | 23.145 | 1.00 | 23.4 | 81500 | 0.58 | 2/2 | |
232524621 | 3.4918 0.0005 | 14.341 | 1.00 | 8.5 | 100000 | 0.65 | 2/9 | hotDA+pair |
1003414425 | 26.46 0.04 | 25.074 | 1.00 | 16.5 | 18089 | 0.173 | 4/4 | |
233177285 | 15.194 0.01 | 4.790 | 1.00 | 6.2 | 6500 | 0.2 | 2/6 | |
232979174 | 8.8045 0.0033 | 8.642 | 1.00 | 5.6 | 12776 | 0.473 | 2/9 | |
1001602651 | 4.9125 0.0018 | 10.704 | 1.00 | 7.2 | 9786 | 0.826 | 5/6 | |
219009693 | 1.5828 0.0004 | 9.973 | 1.00 | 7.3 | 14000 | 0.23 | 4/4 | |
364160098 | 0.19744 6e-05 | 4.307 | 1.00 | 5.7 | 20000 | 0.2 | 1/1 | |
471013541 | 0.549 0.0005 | 4.071 | 1.00 | 5.6 | 88642 | 0.51 | 1/1 | |
951282189 | 2.9618 0.0005 | 95.469 | 0.99 | 110.8 | 19099 | 0.174 | 2/2 | DAB+pair |
951635495 | 23.45 0.03 | 7.002 | 1.00 | 6.6 | 61045 | 0.5 | 2/2 | |
1924161836 | 25.132 0.024 | 14.278 | 1.00 | 12.4 | 71080 | 0.482 | 2/2 | |
410414842 | 14.9 0.4 | 6.417 | 1.00 | 9.7 | 70000 | 0.5 | 1/1 | |
2052284133 | 16.652 0.03 | 12.980 | 1.00 | 14.3 | 34299 | 0.51 | 2/2 | |
251080865 | 32.01 0.06 | 179.265 | 1.00 | 38.0 | 7170 | 0.744 | 2/2 | DCP7 |
164681986 | 3.9437 0.0009 | 33.642 | 1.00 | 31.3 | 8471 | 1.006 | 2/2 | |
392797216 | 0.6808 0.0004 | 6.037 | 1.00 | 6.0 | 6814 | 0.749 | 2/2 | DAH |
157201137 | 0.776229 3.3e-05 | 10.198 | 1.00 | 10.9 | 11254 | 0.486 | 2/2 | CV |
219868627 | 5.9071 0.0017 | 43.787 | 1.00 | 18.3 | 18000 | 0.18 | 12/12 | DAZ+pair |
155872634 | 3.9145 0.0007 | 26.280 | 1.00 | 16.7 | 24589 | 0.697 | 6/6 | |
357389336 | 42.46 0.08 | 7.007 | 1.00 | 6.1 | 21785 | 0.49 | 2/5 | DA+K |
313894558 | 19.05 0.04 | 4.861 | 0.99 | 5.2 | 18610 | 0.61 | 1/2 | |
441569276 | 2.47353 0.00023 | 90.759 | 1.00 | 55.4 | 19690 | 0.43 | 6/6 | DA+M |
[…]

### v2

@@ 5 Conclusion
| 144002497 | 7.6213 0.0022 | 48.042 | 1.00 | 42.8 | 72910 | 0.606 | 2/2 | |
| 837076117 | 11.468 0.007 | 7.899 | 1.00 | 11.2 | 81335 | 0.65 | 1/2 | DA |
| 307982318 | 1.53006 7e-05 | 4.803 | 1.00 | 5.0 | 31880 | 0.67 | 1/3 | hotDAV |
| 88564975 | 8.9702 0.003 | 173.775 | 1.00 | 92.9 | 80000 | 0.67 | 3/3 | |
| 136884288 | 0.6053 0.0006 | 7.730 | 1.00 | 13.6 | 8440 | 0.81 | 1/1 | DAHe |
| 383647517 | 13.924 0.014 | 13.263 | 1.00 | 6.2 | 6754 | 0.545 | 7/16 | DQpec |
| 1102584065 | 26.187 0.022 | 17.910 | 1.00 | 9.1 | 110000 | 0.57 | 12/13 | DO |
| 436747174 | 4.166 0.026 | 290.193 | 1.00 | 59.7 | 22500 | 0.532 | 1/1 | DA-CV+M |
| 434196824 | 3.904 0.023 | 53.066 | 1.00 | 62.4 | 14491 | 0.258 | 1/1 | |
| 345265261 | 3.759 0.022 | 7.601 | 1.00 | 12.2 | 38780 | 0.843 | 1/1 | DA |
| 321272216 | 4.408 0.031 | 34.385 | 0.99 | 40.0 | 12597 | 0.117 | 1/1 | |
| 248861930 | 12.93 0.26 | 23.353 | 1.00 | 28.0 | 36903 | 0.521 | 1/1 | DBA |
| 247992671 | 15.4 0.4 | 8.742 | 1.00 | 10.8 | 8000 | 0.23 | 1/1 | |
| 241256067 | 15.778 0.028 | 5.352 | 1.00 | 6.0 | 4730 | 0.495 | 1/2 | DC |
| 124894242 | 0.32555 0.00017 | 4.075 | 1.00 | 6.9 | 15919 | 0.435 | 1/1 | |
| 154942231 | 4.4236 0.0022 | 7.839 | 1.00 | 6.4 | 12000 | 0.15 | 3/5 | |
| 649932200 | 0.9712 0.0005 | 31.107 | 1.00 | 28.4 | 26461 | 1.036 | 3/3 | |
| 52584399 | 3.801 0.008 | 112.873 | 1.00 | 66.8 | 19505 | 0.249 | 3/3 | |
| 1883129777 | 1.9446 0.0004 | 143.936 | 1.00 | 71.9 | 16480 | 0.229 | 3/3 | |
| 1881564044 | 6.137 0.004 | 6.042 | 1.00 | 7.4 | 34000 | 0.2 | 2/3 | DAH |
| 1101460688 | 23.24 0.03 | 10.905 | 1.00 | 6.9 | 88000 | 0.656 | 2/3 | |
| 22014729 | 7.6477 0.0032 | 135.408 | 1.00 | 55.8 | 36580 | 0.44 | 1/2 | |
| 31798074 | 21.828 0.017 | 42.369 | 1.00 | 5.3 | 12011 | 0.398 | 1/10 | |
| 262921847 | 13.892 0.004 | 52.290 | 1.00 | 20.7 | 7000 | 0.23 | 12/12 | |
| 311886800 | 10.2157 0.0025 | 32.576 | 1.00 | 19.4 | 100000 | 0.6 | 6/6 | |
| 178994042 | 0.2138 7e-05 | 4.400 | 1.00 | 5.6 | 19770 | 0.57 | 1/1 | |
| 220478684 | 16.611 0.010 | 25.143 | 1.00 | 12.6 | 44589 | 0.483 | 19/20 | |
| 381976323 | 8.722 0.01 | 133.547 | 1.00 | 45.2 | 4000 | 0.2 | 11/11 | |
| 235072034 | 84.0 4.0 | 20.540 | 1.00 | 10.5 | 3500 | 0.23 | 2/3 | |
| 219244444 | 7.2886 0.0027 | 245.106 | 1.00 | 12.5 | 6440 | 0.25 | 3/5 | DA7+M+planet |
| 245834374 | 33.6 1.8 | 7.275 | 1.00 | 6.8 | 34210 | 0.69 | 1/1 | DA+M |
| 85146644 | 5.317 0.0011 | 68.977 | 1.00 | 51.4 | 60260 | 0.41 | 3/3 | |
| 107310223 | 11.28 0.23 | 21.505 | 1.00 | 30.9 | 48116 | 0.366 | 1/1 | DA1+dM |
| 1101900758 | 2.65191 0.0003 | 31.335 | 1.00 | 21.4 | 23941 | 0.398 | 5/5 | |
| 160119816 | 21.453 0.017 | 112.794 | 1.00 | 43.3 | 90000 | 0.56 | 9/9 | |
| 300284179 | 12.65 0.29 | 12.426 | 1.00 | 13.8 | 38490 | 0.54 | 1/1 | |
| 126910998 | 19.4 0.6 | 36.753 | 1.00 | 54.9 | 23075 | 0.819 | 1/1 | DA+M |
| 218971976 | 6.7081 0.001 | 263.431 | 1.00 | 65.1 | 39260 | 0.38 | 2/2 | DA+M |
| 321979116 | 2.69506 0.00016 | 92.402 | 1.00 | 36.5 | 8650 | 0.96 | 2/2 | DAP5.8 |
| 183533908 | 2.55267 0.00024 | 11.122 | 1.00 | 13.6 | 7638 | 0.74 | 1/1 | |
| 630539472 | 1.592 0.004 | 4.231 | 1.00 | 7.4 | 10027 | 0.821 | 1/1 | |
| 201255204 | 6.34 0.03 | 59.895 | 1.00 | 52.0 | 4000 | 0.2 | 2/2 | DBH |
| 439917321 | 12.98 0.28 | 61.269 | 1.00 | 48.6 | 19193 | 0.57 | 1/1 | DA+M |
| 55096188 | 2.75096 0.00021 | 4.379 | 1.00 | 6.5 | 10000 | 0.6 | 2/3 | |
| 53190694 | 9.923 0.004 | 12.529 | 1.00 | 9.3 | 30502 | 0.41 | 3/4 | DA3+dM |
| 952264940 | 1.78894 0.00018 | 7.713 | 1.00 | 8.6 | 25692 | 1.151 | 1/2 | DBH |
| 219769847 | 8.72 0.13 | 13.789 | 1.00 | 20.7 | 21620 | 0.72 | 1/1 | DA3+dM |
| 292114384 | 0.29255 0.00014 | 4.110 | 1.00 | 5.1 | 18943 | 0.57 | 1/1 | |
| 66714896 | 12.579 0.009 | 10.848 | 1.00 | 8.8 | 63073 | 0.69 | 1/3 | |
| 349233389 | 52.0 5.0 | 31.595 | 1.00 | 25.7 | 23690 | 0.58 | 1/1 | DA2+M |
| 335501964 | 6.7654 0.0025 | 5.380 | 1.00 | 8.1 | 90000 | 0.53 | 1/2 | |
| 165916724 | 0.092325 5e-07 | 4.156 | 1.00 | 5.2 | 9557 | 0.883 | 1/3 | DAH |
| 274039489 | 7.32 0.09 | 67.842 | 1.00 | 58.5 | 42370 | 0.66 | 1/1 | DA1.2+pair |
| 328029653 | 14.157 0.004 | 19.335 | 0.99 | 12.5 | 6340 | 0.53 | 5/5 | |
| 311920919 | 0.1614889 1e-06 | 4.611 | 1.00 | 6.2 | 14116 | 0.68 | 1/4 | |
| 155871645 | 0.1385659 8e-07 | 6.552 | 1.00 | 5.5 | 34050 | 0.77 | 3/5 | |
| 344927121 | 2.53418 0.00018 | 113.339 | 1.00 | 67.3 | 26088 | 0.381 | 4/4 | |
| 148719841 | 8.549 0.004 | 17.920 | 1.00 | 18.7 | 6169 | 0.72 | 2/2 | |
| 415714190 | 0.384926 4e-06 | 9.333 | 1.00 | 8.4 | 7690 | 0.636 | 3/3 | |
| 277713491 | 8.9986 0.0032 | 11.795 | 1.00 | 9.2 | 7518 | 0.578 | 3/4 | |
| 471013515 | 6.05 0.07 | 27.262 | 1.00 | 34.8 | 55000 | 0.6 | 1/1 | |
| 5393020 | 1.5786 0.00014 | 54.426 | 1.00 | 49.4 | 8000 | 0.35 | 1/2 | |
| 610754879 | 0.91146 9e-05 | 11.175 | 1.00 | 10.9 | 47617 | 1.341 | 2/3 | |
| 365207316 | 2.539 0.005 | 6.316 | 1.00 | 7.1 | 10520 | 0.226 | 2/2 | |
| 425079955 | 17.076 0.016 | 20.821 | 1.00 | 16.2 | 14252 | 0.285 | 2/2 | |
| 1271382181 | 7.426 0.005 | 40.083 | 1.00 | 21.8 | 24015 | 0.25 | 8/8 | |
| 115970839 | 2.6994 0.0004 | 41.932 | 1.00 | 30.0 | 8825 | 0.219 | 5/5 | DA+M: |
| 159819681 | 10.728 0.006 | 29.967 | 1.00 | 22.8 | 27876 | 0.573 | 5/5 | |
| 158971558 | 6.327 0.004 | 7.300 | 1.00 | 7.3 | 14319 | 0.323 | 2/2 | |
| 1550807820 | 0.566897 1.6e-05 | 5.636 | 1.00 | 8.6 | 44523 | 1.293 | 5/5 | |
| 7622101 | 8.5708 0.0028 | 11.742 | 1.00 | 7.6 | 14159 | 0.2 | 7/7 | |
| 349993999 | 4.381 0.0005 | 7.642 | 1.00 | 7.0 | 24751 | 0.756 | 1/1 | |
| 1688534307 | 0.0902002 4e-07 | 4.253 | 1.00 | 5.3 | 66522 | 0.637 | 1/2 | |
| 62918265 | 4.6612 0.0008 | 12.845 | 0.99 | 13.1 | 8133 | 0.724 | 6/6 | |
| 1922433415 | 2.03387 0.00015 | 129.416 | 1.00 | 60.4 | 21155 | 0.214 | 2/2 | |
| 100469915 | 12.936 0.006 | 18.310 | 1.00 | 21.6 | 12644 | 0.152 | 2/2 | |
| 2026050241 | 8.0844 0.0024 | 6.393 | 1.00 | 5.7 | 48831 | 0.489 | 1/2 | |
| 60112260 | 7.37 0.1 | 14.436 | 1.00 | 20.0 | 7682 | 0.756 | 1/1 | |
| 269937578 | 7.2969 0.0019 | 79.671 | 1.00 | 45.4 | 32755 | 0.274 | 4/5 | |
| 2052430471 | 0.125404 2.9e-05 | 4.277 | 1.00 | 5.4 | 28678 | 0.368 | 1/1 | |
| 124573902 | 7.9918 0.0035 | 21.546 | 1.00 | 21.9 | 12173 | 0.371 | 2/2 | wd+pair |
| 270432557 | 9.311 0.01 | 6.316 | 1.00 | 5.4 | 10000 | 0.23 | 2/3 | |
| 43529091 | 9.44 0.14 | 11.479 | 1.00 | 14.0 | 11006 | 0.17 | 1/1 | |
| 278861557 | 3.1624 0.0004 | 26.520 | 1.00 | 7.8 | 10460 | 0.23 | 19/20 | wd+pair |
| 62846288 | 0.500906 1.4e-05 | 23.831 | 1.00 | 17.9 | 16268 | 0.935 | 3/4 | |
| 56813164 | 8.05 0.11 | 25.988 | 1.00 | 34.7 | 17304 | 0.39 | 1/1 | DA+M |
| 953086708 | 0.088138 1.3e-05 | 5.293 | 1.00 | 7.4 | 7735 | 0.564 | 1/1 | DAEH |
| 95475692 | 1.94319 0.00021 | 45.117 | 1.00 | 36.2 | 13737 | 0.428 | 2/2 | DA+M |
| 68015843 | 3.015 0.0018 | 21.609 | 1.00 | 21.7 | 12392 | 0.4 | 2/2 | DA+M4.5 |
| 219485299 | 1.9770 0.0020 | 4.513 | 1.00 | 18.9 | 10307 | 0.63 | 3/3 | DA+M |
| 1507548590 | 1.9272 0.0004 | 50.124 | 1.00 | 49.2 | 25089 | 0.214 | 2/2 | |
| 1101327387 | 1.30545 9e-05 | 21.739 | 1.00 | 19.5 | 15659 | 1.13 | 2/2 | DAH |
| 416538823 | 16.021 0.011 | 335.044 | 1.00 | 26.1 | 18840 | 0.8 | 4/4 | DA+M |
| 7983187 | 14.32 0.16 | 77.693 | 1.00 | 45.2 | 17210 | 0.59 | 2/2 | DA3+dM |
| 251857373 | 2.31675 0.00012 | 175.807 | 1.00 | 110.6 | 40698 | 0.56 | 2/2 | |
| 408015814 | 29.52 0.05 | 13.089 | 1.00 | 11.8 | 76010 | 0.061 | 2/2 | |
| 200725303 | 9.48 0.07 | 35.024 | 1.00 | 33.5 | 90000 | 0.75 | 2/2 | |
| 159894126 | 0.46415 0.00017 | 4.410 | 1.00 | 5.9 | 8627 | 0.613 | 1/2 | |
| 651923454 | 42.4 2.9 | 16.656 | 1.00 | 11.9 | 21644 | 0.746 | 1/1 | wd+pair |
| 651337335 | 30.5 1.5 | 6.331 | 1.00 | 9.7 | 21796 | 0.217 | 1/1 | |
| 650999219 | 24.6 0.5 | 24.641 | 1.00 | 26.3 | 70000 | 0.49 | 2/2 | |
| 93031595 | 9.94 0.16 | 17.306 | 1.00 | 25.7 | 8751 | 0.911 | 1/1 | DZ |
| 685012668 | 2.8735 0.0018 | 107.688 | 1.00 | 44.9 | 20041 | 0.238 | 5/5 | |
| 672383331 | 17.8 0.5 | 7.107 | 1.00 | 10.9 | 21076 | 0.279 | 1/1 | |
| 257689590 | 13.35 0.13 | 5.132 | 0.99 | 5.2 | 8629 | 0.183 | 2/2 | |
| 455072135 | 11.36 0.21 | 90.850 | 1.00 | 49.0 | 12965 | 0.116 | 1/1 | EB |
| 941411508 | 10.547 0.006 | 8.630 | 1.00 | 9.1 | 40000 | 0.556 | 2/2 | |
| 942842236 | 8.233 0.004 | 10.210 | 1.00 | 10.1 | 77164 | 0.75 | 2/2 | |
| 332828037 | 2.07518 0.00024 | 15.973 | 1.00 | 17.4 | 13936 | 0.382 | 2/2 | |
| 939953390 | 29.85 0.05 | 9.949 | 1.00 | 6.7 | 63208 | 0.667 | 3/3 | |
| 952000432 | 21.158 0.025 | 23.145 | 1.00 | 23.4 | 81500 | 0.58 | 2/2 | |
| 232524621 | 3.4918 0.0005 | 14.341 | 1.00 | 8.5 | 100000 | 0.65 | 2/9 | hotDA+pair |
| 1003414425 | 26.46 0.04 | 25.074 | 1.00 | 16.5 | 18089 | 0.173 | 4/4 | |
| 233177285 | 15.194 0.01 | 4.790 | 1.00 | 6.2 | 6500 | 0.2 | 2/6 | |
| 232979174 | 8.8045 0.0033 | 8.642 | 1.00 | 5.6 | 12776 | 0.473 | 2/9 | |
| 1001602651 | 4.9125 0.0018 | 10.704 | 1.00 | 7.2 | 9786 | 0.826 | 5/6 | |
| 219009693 | 1.5828 0.0004 | 9.973 | 1.00 | 7.3 | 14000 | 0.23 | 4/4 | |
| 364160098 | 0.19744 6e-05 | 4.307 | 1.00 | 5.7 | 20000 | 0.2 | 1/1 | |
| 471013541 | 0.549 0.0005 | 4.071 | 1.00 | 5.6 | 88642 | 0.51 | 1/1 | |
| 951282189 | 2.9618 0.0005 | 95.469 | 0.99 | 110.8 | 19099 | 0.174 | 2/2 | DAB+pair |
| 951635495 | 23.45 0.03 | 7.002 | 1.00 | 6.6 | 61045 | 0.5 | 2/2 | |
| 1924161836 | 25.132 0.024 | 14.278 | 1.00 | 12.4 | 71080 | 0.482 | 2/2 | |
| 410414842 | 14.9 0.4 | 6.417 | 1.00 | 9.7 | 70000 | 0.5 | 1/1 | |
| 2052284133 | 16.652 0.03 | 12.980 | 1.00 | 14.3 | 34299 | 0.51 | 2/2 | |
| 251080865 | 32.01 0.06 | 179.265 | 1.00 | 38.0 | 7170 | 0.744 | 2/2 | DCP7 |
| 164681986 | 3.9437 0.0009 | 33.642 | 1.00 | 31.3 | 8471 | 1.006 | 2/2 | |
| 392797216 | 0.6808 0.0004 | 6.037 | 1.00 | 6.0 | 6814 | 0.749 | 2/2 | DAH |
| 157201137 | 0.776229 3.3e-05 | 10.198 | 1.00 | 10.9 | 11254 | 0.486 | 2/2 | CV |
| 219868627 | 5.9071 0.0017 | 43.787 | 1.00 | 18.3 | 18000 | 0.18 | 12/12 | DAZ+pair |
| 155872634 | 3.9145 0.0007 | 26.280 | 1.00 | 16.7 | 24589 | 0.697 | 6/6 | |
| 357389336 | 42.46 0.08 | 7.007 | 1.00 | 6.1 | 21785 | 0.49 | 2/5 | DA+K |
| 313894558 | 19.05 0.04 | 4.861 | 0.99 | 5.2 | 18610 | 0.61 | 1/2 | |
| 441569276 | 2.47353 0.00023 | 90.759 | 1.00 | 55.4 | 19690 | 0.43 | 6/6 | DA+M |
[…]